Feed aggregator

HYST AREA 6-2-4 Cycles Magnets

Recent Log Entries -

Log Number: 4097704Logbook(s): Body: 

HYST AREA Version 6-2-4 has successfully cycled the following:

Area.Halls.HallC

The following magnets were REMOVED from the CYCLING LIST by the user:
MBC3C01H
MBC3C02V
MBD3C00AV
MBD3C00V
MQA3C01
Total Time: 5.733 mins

Operator SIGN IN

Recent Log Entries -

Log Number: 4097700Logbook(s): Body:  ACCELERATOR OPERATOR SIGN IN FORM OPERATOR:    Deir
DATE:             12/07/2022
SHIFT:               Owl

The operator has read the following:
  • ELOG
  • PD Shift Plan
  • Radiation Surveys for PSS Segments in Controlled Access
  • PSS Log

Beam recovery to Hall A during swing shift in 12/06

Recent Log Entries -

Log Number: 4097698Logbook(s): Entry Makers: jpoudelBody: 

Beam tuning to Hall A started after 15 pm and first set of harp scan has sigma-Y to target less than 0.30 microns (beam operation restriction states both x and y sigma >0.3), so we asked MCC to increase the spot size in Y. Second harp scan didn't improve Y, but X was better (~0.38). MCC performed some more harp scans to make both X and Y >0.32 (as prescribed by Simona during the start-up of GEnII). Harp scan around 16:45 was much better (0.37,0.32).

Pages

Subscribe to cebaf2013 aggregator